# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ 5.2.2 Token-level ablation
In addition, removing template tokens causes the LLMs to completely lose their ability to solve tasks via ICL with an overall task accuracy performance of 0% for all sizes and all tasks. We hypothesize that this is because the model no longer has an explicit cue to generate the target label, which is further discussed in Section 6. In this case, if we add back the last prompt token after the next-line token, the results return to their original level due to the introduction of a template token. This finding confirms previous claims that preserving the format of ICL prompts plays a significant role in retaining the task performance (Min et al., 2022).
Roles of different types of tokens
To summarize, template and stopword tokens are the most performance-critical. Template token representations significantly improve the task performance, while the representations of stopword tokens play a more supportive role in the spectrum of performance-critical tokens by summarizing the information of content tokens. In contrast, content token representations do not directly contribute to the performance but instead indirectly provide information that is aggregated into the other two types of tokens. We discuss possible applications of these findings in Appendix Q. This finding further raises an additional question: What are the characteristics for a token to be perceived by a LLM as performance-critical?
6 Performance-Critical Token Analyses
To answer the above question, we provide analyses of the tokens whose representations we believe mainly store information that directly affects the performance of a task drastically. We focus on the template tokens since, as evidenced by the findings in Section 5.2.1 and 5.2.2, their representations are the most important to maintaining task performance. Our analysis focuses on the distinguishing characteristics of performance-critical tokens, while we also examine the effects of each part of template tokens in Appendix C.
@@ 6 Performance-Critical Token Analyses
We design several experiments to test if these characteristics affect the impact of performance-critical tokens, by disrupting each characteristic in the ICL prompts. A characteristic is related if there is a performance drop after the disruption. The disruption is achieved by replacing the template tokens with different kinds of random string templates, shown in Table 5. We use 5 different random string templates which are attached to Appendix T and average all the results for each setting.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
Open Llama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
Swap | 64.4 | 86.8 | 21.7 | 58.7 | 60.6 | 54.6 | 57.8 | |
Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
Swap | 70.2 | 11.4 | 44.3 | 58.2 | 64.5 | 50.1 | 49.8 | |
Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
Swap | 81.5 | 67.4 | 36.4 | 75.9 | 69.1 | 52.1 | 63.7 | |
Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
Swap | 84.5 | 94.9 | 60.8 | 75.5 | 68.0 | 55.5 | 73.2 | |
Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 |
Lexical meaning
A performance-critical token might be more impactful on the performance based on its lexical meaning. One possible hypothesis is that if the token carries specific task-related meanings like “Answer”, it is more likely to serve as a performance-critical token.
@@
Repetition
The impact of performance-critical tokens could also be influenced by their repetition throughout the prompt. Intuitively, via the attention mechanism, repetitive patterns are more likely to propagate information through the processing of text. Yan et al. (2024) propose self-reinforcement in in-context learning, also suggesting that repetition could be a significant factor in ICL.
We explore the repetition characteristic by comparing the results of the previously discussed Randomfixed experiment with an experiment replacing and with different random strings (Randomnonfixed), thus breaking the repetition of template tokens present in ICL demonstrations. We further conduct supplemental experiments using template tokens with specific lexical meanings for comparison, as detailed in Appendix U.
Table 7 shows that without consistent repetition of performance-critical tokens, performance decreases for most models. This suggests that necessary information may not have been properly accumulated in the template token representations. These experiments demonstrate that repetition significantly influences the impact of performance-critical tokens. The results align with previous findings, reinforcing our claim that repetition is a key characteristic of performance-critical tokens.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 |
Randomnonfixed | 30.2 | 71.4 | 17.1 | 18.6 | 47.9 | 47.7 | 38.8 | |
Llama 7B | Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 |
Randomnonfixed | 15.5 | 11.6 | 10.4 | 1.8 | 4.6 | 25.6 | 11.6 | |
Llama 13B | Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 |
Randomnonfixed | 32.1 | 34.5 | 19.2 | 6.0 | 21.0 | 32.8 | 24.3 | |
Llama 33B | Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 |
@@
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 70.7 | 51.7 | 40.4 | 53.5 | 50.2 | 48.6 | 53.3 |
Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 | |
Llama 7B | Standard ICL | 72.3 | 77.4 | 54.1 | 64.7 | 53.0 | 64.4 | 64.3 |
Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 | |
Llama 13B | Standard ICL | 82.0 | 72.0 | 60.1 | 75.9 | 60.4 | 18.8 | 70.1 |
Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 | |
Llama 33B | Standard ICL | 85.3 | 88.3 | 71.2 | 75.5 | 64.1 | 45.5 | 76.9 |
@@
Structural cues
Beyond lexical meaning and repetition, the influence of performance-critical tokens may also depend on how ICL prompts are formatted. ICL prompts often include structural cues to assist the model to differentiate between elements with distinct roles, such as task inputs and target labels, within a demonstration. For example, template tokens (i.e., and ) delimit demonstration examples and labels, while stopword tokens (e.g., ,”, .”, “:”, etc.) structure content words into sentence components. Examples of how performance-critical tokens delimit ICL prompts are shown in Appendix W. These structural cues are similar to those in LLM pretraining data (e.g., column names in SQL tables), suggesting that pretraining on such data enables the model to recognize the structuring role of performance-critical tokens, allowing their representations to store higher-level information.
To assess the structuring characteristic of performance-critical tokens, we perturb the structure of one-shot ICL prompts in two stages, where the one-shot setting could eliminate repetition as a confounding factor. First, we disrupt the lexical meaning of template tokens, since token meaning helps LLMs distinguishing the different parts of a prompt. Then, we remove all template tokens to eliminate any source of structure cues. Table 8 shows that disrupting structural cues decreases performance, highlighting their importance. Consistent with Section 5.2.2, removing all template tokens results in 0% performance due to the complete elimination of structural cues. Supplemental experiments in Appendix V further support this from a representation-level ablation perspective.
@@ 7 Conclusion
Limitations
In this paper, the token categorization is performed manually, leaving room for further refinement, leaving the exploration of other specific content tokens as performance-critical tokens in certain contexts to future work. While the results provide robust support to our categorization, the identification process itself lacks precision. For instance, stopwords may only represent a subset of all in-context performance-critical tokens. The manual nature of our categorization limits our ability to comprehensively track these tokens. Moreover, our experiments are limited to classification, machine translation, and question answewring datasets, suggesting that our conclusions should be further validated for other tasks. Additionally, our focus on performance-critical tokens, whose representations could impact task performance, may overlook other tokens responsible for other possible functions. Another limitation of our study is that we focus exclusively on in-context learning scenarios, meaning that our findings may not be directly applicable to zero-shot learning scenarios.
@@ Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we have made several efforts that are documented throughout this paper. Our experiments utilize the open-source models described in Section 4. The prompts and templates used in our experiments are detailed in Section 3.1, Section 6 of the main text and in Appendix D, Appendix E, Appendix K, Appendix T, Appendix U. The stopword token list used in our experiments is shown in Appendix H. The complete code for our implementation, including all inference processes, is provided in the supplementary materials. We employed random seeds ranging from 1 to 15 to ensure consistent results across experiments, as specified in Section 4 and the supplementary code. All datasets used in our experiments are described comprehensively in Section 4, and the supplementary code includes all data processing steps and any preprocessing applied. We encourage other researchers to consult these references for replicating our findings.
Datasets | Notations | Examples |
@@
Models | Settings | RTE | Settings | RTE | ||
with “:” | w/o “:” | with “:” | w/o “:” | |||
Llama 7B | temp with | 66.3 | 59.5 | temp w/o | 40.7 | 42.5 |
58.9 | 56.5 | 43.7 | 49.9 | |||
56.7 | 56.7 | 56.0 | 55.6 | |||
Llama 13B | temp with | 68.5 | 59.8 | temp w/o | 57.5 | 53.7 |
65.5 | 59.0 | 53.6 | 52.8 | |||
61.2 | 58.4 | 54.8 | 53.7 | |||
Llama 33B | temp with | 79.0 | 77.1 | temp w/o | 71.8 | 65.8 |
77.4 | 75.4 | 70.6 | 67.8 | |||
72.8 | 70.0 | 67.0 | 61.3 |
Appendix B More Related Work
B.1 Working mechanisms of ICL
Since the proposal of in-context learning (Brown et al., 2020), its working mechanisms have been extensively studied by the research community (Min et al., 2022; Liu et al., 2021; Olsson et al., 2022; Bhattamishra et al., 2023). Min et al. (2022) suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, Yoo et al. (2022) propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. Xie et al. (2021) explain ICL as implicit Bayesian inference, while Akyürek et al. (2022) explore ICL learning process using linear models. Theoretical explanations (Guo et al., 2023; Bai et al., 2023; Li et al., 2023b) and gradient descent explanations have also been proposed. Mao et al. (2024) analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (Han et al., 2023). Zhao et al. (2024) propose to use coordinate systems to understand the working mechanism of in-context learning. Zhou et al. (2023) propose a comprehensive survey on the interpretation and analysis of in-context learning.
Additional analyses exploring different aspects of ICL have also been studied. For instance, order sensitivity where task performance fluctuates based on the order of the same ICL demonstrations has been identified as a limitation of ICL (Lu et al., 2022). Yan et al. (2024) propose that repetitive patterns in the prompt could affect the ICL performance in both positive and negative ways. Pan et al. (2023) analyze the ICL process by disentangling it into task recognition and task learning. Madaan and Yazdanbakhsh (2022) propose to define text and patterns while using counterfactual prompting for attributing token importance in chain-of-thought techniques.
@@ B.1 Working mechanisms of ICL
B.2 Function vectors of in-context learning
Todd et al. (2023) and Hendel et al. (2023) provide evidence of function vectors that store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt. These vectors can then be added to, or used to replace, the corresponding vectors in a zero-shot example, achieving results comparable to those obtained when the model uses all demonstrations as context. In addition, Liu et al. (2023a) also propose using an in-context vector to represent the target task and applying feature shifting to query examples. They first feed each input and its corresponding target separately into an LLM, then concatenate all the latent states. A PCA method is applied to derive a vector that is more closely aligned with the task. Finally, Wang et al. (2023) propose that label words in the demonstration examples function as information anchors by aggregating the information from previous demonstrations and providing it to the test example. This finding suggests that we may view label tokens as satisfying our definition of performance-critical tokens.

### v4

@@ 5.2 Ablation on token types
In addition, removing template tokens causes the LLMs to completely lose their ability to solve tasks via ICL with an overall task accuracy performance of 0% for all sizes and all tasks. We hypothesize that this is because the model no longer has an explicit cue to generate the target label, which is further discussed in Section 8. In this case, if we add back the last prompt token after the next-line token, the results return to their original level due to the introduction of a template token. This finding confirms previous claims that the format of ICL prompts plays a significant role in retaining performance (min2022rethinking).
Roles of different types of tokens. To summarize, template and stopword tokens are the most performance-critical. Template token representations significantly improve the task performance, while the representations of stopword tokens play a more supportive role in the spectrum of performance-critical tokens by summarizing the information of content tokens. In contrast, content token representations do not directly contribute to the performance but instead indirectly provide information that is aggregated into the other two types of tokens. We discuss possible applications of these findings in Appendix R. This finding further raises an additional question: What are the characteristics for a token to be perceived by a LLM as performance-critical?
@@
6 Performance-Critical Token Analyses
To answer the above question, we provide analyses of the tokens whose representations we believe mainly store information that directly affects the performance of a task drastically. We focus on the template tokens since, as evidenced by the findings in Section 5.2, their representations are the most important to maintaining task performance. Our analysis focuses on the distinguishing characteristics of performance-critical tokens, while we also examine the effects of each part of template tokens in Appendix D.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | |||
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 | |||
| Swap | 64.4 | 86.8 | 21.7 | 58.7 | 60.6 | 54.6 | 57.8 | ||||
| Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 | ||||
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 | |||
| Swap | 70.2 | 11.4 | 44.3 | 58.2 | 64.5 | 50.1 | 49.8 | ||||
| Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 | ||||
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 | |||
| Swap | 81.5 | 67.4 | 36.4 | 75.9 | 69.1 | 52.1 | 63.7 | ||||
| Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 | ||||
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 | |||
| Swap | 84.5 | 94.9 | 60.8 | 75.5 | 68.0 | 55.5 | 73.2 | ||||
| Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 |
@@ 6 Performance-Critical Token Analyses
We design several experiments to test if these characteristics affect the impact of performance-critical tokens, by disrupting each characteristic in the ICL prompts. A characteristic is related if there is a performance drop after the disruption. The disruption is achieved by replacing the template tokens with different kinds of random string templates, shown in Table 6. We use 5 different random string templates which are attached to Appendix U and average all the results for each setting.
Lexical meaning. A performance-critical token might be more impactful on the performance based on its lexical meaning. One hypothesis is that if the token carries task-related meanings like “Answer”, it is more likely to serve as a performance-critical token.
@@
Repetition. The impact of performance-critical tokens could also be influenced by their repetition throughout the prompt. Intuitively, via the attention mechanism, repetitive patterns are more likely to propagate information through the processing of text. yan2023understanding propose self-reinforcement in in-context learning, also suggesting that repetition could be a significant factor in ICL.
We explore the repetition characteristic by comparing the results of the previously discussed Randomfixed experiment with an experiment replacing and with different random strings (Randomnonfixed), thus breaking the repetition of template tokens present in ICL demonstrations. We further conduct experiments using template tokens with specific lexical meanings for comparison, as detailed in Appendix V.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Randomfixed | 57.5 | 71.4 | 32.4 | 51.2 | 53.3 | 49.8 | 52.6 | ||
| Randomnonfixed | 30.2 | 71.4 | 17.1 | 18.6 | 47.9 | 47.7 | 38.8 | |||
Randomfixed | 19.5 | 11.4 | 13.2 | 7.4 | 19.7 | 21.7 | 15.5 | ||
| Randomnonfixed | 15.5 | 11.6 | 10.4 | 1.8 | 4.6 | 25.6 | 11.6 | |||
Randomfixed | 52.1 | 76.8 | 27.7 | 48.9 | 55.7 | 34.5 | 49.3 | ||
| Randomnonfixed | 32.1 | 34.5 | 19.2 | 6.0 | 21.0 | 32.8 | 24.3 | |||
Randomfixed | 78.7 | 92.5 | 52.2 | 75.8 | 68.9 | 41.1 | 68.2 | ||
@@
Table 8 shows that without consistent repetition of performance-critical tokens, performance decreases for most models. This suggests that necessary information may not have been properly accumulated in the template token representations. These experiments demonstrate that repetition significantly influences the impact of performance-critical tokens. The results align with previous findings, reinforcing our claim that repetition is a key characteristic.
Structural cues. Beyond lexical meaning and repetition, the influence of performance-critical tokens may also depend on how ICL prompts are formatted. ICL prompts often include structural cues to assist the model to differentiate between elements with distinct roles, such as task inputs and target labels, within a demonstration. For example, template tokens (i.e., and ) delimit demonstration examples and labels, while stopword tokens (e.g., ,”, .”, “:”, etc.) structure content words into sentence components. Examples of how performance-critical tokens delimit ICL prompts are shown in Appendix X. These structural cues are similar to those in LLM pretraining data (e.g., column names in SQL tables), suggesting that pretraining on such data enables the model to recognize the structuring role of performance-critical tokens, allowing the representations to store higher-level information.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Standard ICL | 70.7 | 51.7 | 40.4 | 53.5 | 50.2 | 48.6 | 53.3 | ||
| Randomfixed | 47.5 | 51.8 | 32.6 | 19.4 | 51.8 | 42.4 | 40.9 | |||
Standard ICL | 72.3 | 77.4 | 54.1 | 64.7 | 53.0 | 64.4 | 64.3 | ||
| Randomfixed | 3.9 | 16.9 | 3.5 | 9.6 | 16.9 | 10.4 | 10.2 | |||
Standard ICL | 82.0 | 72.0 | 60.1 | 75.9 | 60.4 | 18.8 | 70.1 | ||
| Randomfixed | 46.1 | 47.5 | 25.0 | 50.8 | 47.5 | 21.4 | 39.7 | |||
Standard ICL | 85.3 | 88.3 | 71.2 | 75.5 | 64.1 | 45.5 | 76.9 | ||
@@
To assess the structuring characteristic of performance-critical tokens, we perturb the structure of one-shot ICL prompts in two stages, where the one-shot setting could eliminate repetition as a confounding factor. First, we disrupt the lexical meaning of template tokens, since token meaning helps LLMs distinguishing the different parts of a prompt. Then, we remove all template tokens to eliminate any source of structure cues. Table 9 shows that disrupting structural cues decreases performance, highlighting their importance. Consistent with Section 5.2, removing all template tokens results in 0% performance due to the complete elimination of structural cues. Supplemental experiments in Appendix W further support this from a representation-level perspective.
@@ 7 Conclusion
Acknowledgements
The authors thank all the reviewers for their suggestions and comments. This work is supported by National Natural Science Foundation of China (No. U21B2009). Jackie Chi Kit Cheung is supported by Canada CIFAR AI Chair program.
@@ Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we have made several efforts that are documented throughout this paper. Our experiments utilize the open-source models described in Section 4. The prompts and templates used in our experiments are detailed in Section 3, Section 6 of the main text and in Appendix E, Appendix F, Appendix L, Appendix U, Appendix V. The stopword token list used in our experiments is shown in Appendix I. The complete code for our implementation, including all inference processes, is provided in the supplementary materials. We employed random seeds ranging from 1 to 15 to ensure consistent results across experiments, as specified in Section 4 and the supplementary code. All datasets used in our experiments are described comprehensively in Section 4, and the supplementary code includes all data processing steps and any preprocessing applied. We encourage other researchers to consult these references for replicating our findings.
Appendix B Limitations
In this paper, the token categorization is performed manually, leaving room for further refinement and the exploration of other specific content tokens as performance-critical tokens in certain contexts to future work. While the results provide robust support to our categorization, the identification process itself lacks precision. For instance, stopwords may only represent a subset of all in-context performance-critical tokens. The manual nature of our categorization limits our ability to comprehensively track these tokens.
Moreover, our experiments are limited to classification, machine translation, and question answewring datasets, suggesting that our conclusions should be further validated for other tasks. Additionally, our focus on performance-critical tokens, whose representations could impact task performance, may overlook other tokens responsible for other possible functions. Another limitation of our study is that we focus exclusively on in-context learning scenarios, meaning that our findings may not be directly applicable to zero-shot learning scenarios.
| Datasets | Notations | Examples | |
@@ Appendix B Limitations
| Models | Settings | RTE | Settings | RTE | ||||
| with “:” | w/o “:” | with “:” | w/o “:” | |||||
temp with | 66.3 | 59.5 | temp w/o | 40.7 | 42.5 | ||
| 58.9 | 56.5 | 43.7 | 49.9 | |||||
| 56.7 | 56.7 | 56.0 | 55.6 | |||||
temp with | 68.5 | 59.8 | temp w/o | 57.5 | 53.7 | ||
| 65.5 | 59.0 | 53.6 | 52.8 | |||||
| 61.2 | 58.4 | 54.8 | 53.7 | |||||
temp with | 79.0 | 77.1 | temp w/o | 71.8 | 65.8 | ||
| 77.4 | 75.4 | 70.6 | 67.8 | |||||
| 72.8 | 70.0 | 67.0 | 61.3 | |||||
Appendix C More Related Work
C.1 Working mechanisms of ICL
Since the proposal of in-context learning (brown2020language), its working mechanisms have been extensively studied by the research community (min2022rethinking; liu2021makes; bhattamishra2023understanding). min2022rethinking suggest that demonstrations primarily provide the label space, the distribution of the input text, and the format of the sequence for the test example. They argue that the precise ground truth labels do not have significant importance. In contrast, yoo2022groundtruth propose a differing view, stating that the impact of the ground truth labels depends on the experimental configuration. xie2021explanation explain ICL as implicit Bayesian inference, while akyrek2022learning explore ICL learning process using linear models. Theoretical explanations (guo2023transformers; bai2023transformers; pmlr-v202-li23l) and gradient descent explanations have also been proposed. mao2024data analyze in-context learning from the perspective of data generation. The perspective of supported training data is also leveraged to analyze ICL (han2023understanding). zhao2024unveiling propose to use coordinate systems to understand the working mechanism of in-context learning. zhou2023mystery propose a comprehensive survey on the interpretation and analysis of in-context learning.
Additional analyses exploring different aspects of ICL have also been studied. For instance, order sensitivity where task performance fluctuates based on the order of the same ICL demonstrations has been identified as a limitation of ICL (lu-etal-2022-fantastically). yan2023understanding propose that repetitive patterns in the prompt could affect the ICL performance in both positive and negative ways. DBLP:conf/acl/PanG0C23 analyze the ICL process by disentangling it into task recognition and task learning. madaan2022text propose to define text and patterns while using counterfactual prompting for attributing token importance in chain-of-thought techniques.
@@ C.1 Working mechanisms of ICL
C.2 Function vectors of in-context learning
todd2023function and hendel2023incontext provide evidence of function vectors that store information used to solve a task in ICL. They probe and extract the hidden representations of the final tokens in the prompt. These vectors can then be added to, or used to replace, the corresponding vectors in a zero-shot example, achieving results comparable to those obtained when the model uses all demonstrations as context. In addition, liu2023incontext also propose using an in-context vector to represent the target task and applying feature shifting to query examples. They first feed each input and its corresponding target separately into an LLM, then concatenate all the latent states. A PCA method is applied to derive a vector that is more closely aligned with the task. Finally, wang2023label propose that label words in the demonstration examples function as information anchors by aggregating the information from previous demonstrations and providing it to the test example. This finding suggests that we may view label tokens as satisfying our definition of performance-critical tokens.